### v2

# Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics

Abstract
Introducing superconductivity into group IV elements by doping has long promised a pathway to introduce quantum functionalities into well-established semiconductor technologies. The non-equilibrium hyperdoping of group III atoms into Si or Ge has successfully shown superconductivity can be achieved, however, the origin of superconductivity has been obscured by structural disorder and dopant clustering. Here, we report the epitaxial growth of hyperdoped Ga:Ge films by molecular beam epitaxy with extreme hole concentrations ($n_{\textup{h}}=4.15\times 10^{21}$ cm-3,  17.9% Ga substitution) that yield superconductivity with a critical temperature of $T_{\textup{c}}=3.5$ K and an out-of-plane critical field of 1 T at 270 mK. Synchrotron-based X-ray absorption and scattering methods reveal that Ga dopants are substitutionally incorporated within the Ge lattice, introducing a tetragonal distortion to the crystal unit cell. Our findings, corroborated by first-principles calculations, suggest that the structural order of Ga dopants creates a narrow band for the emergence of superconductivity in Ge, establishing hyperdoped Ga:Ge as a low-disorder, epitaxial superconductor-semiconductor platform.
preprint: APS/123-QED
Although the realization of superconductivity in group IV elements has been previously demonstrated [1, 2, 3, 4], difficulties such as dopant segregation and precipitation, nanocrystal formation, and incoherent interfaces impede their integration into a new generation of quantum devices. As a prototypical semiconductor and foundry-ready material, germanium is a key component of devices that combine macroscopic superconducting coherence with the microscopic degrees of freedom in semiconductors [5, 6, 7, 8]. While epitaxial superconductor-semiconductor (S-Sm) interfaces have been demonstrated with III-V semiconductors [9, 10], realizing full S-Sm-S heterostructures remains challenging and has limited their applicability in quantum devices [11, 12]. This difficulty has renewed interest in doping-based superconductivity in group IV materials, a concept explored theoretically by Cohen in the 1960s [13], as recent efforts to induce superconductivity through proximity effects [7, 8] and alloying [14] have met with mixed results. This route sidesteps disordered interfaces between dissimilar materials by inducing superconductivity within an epitaxial semiconductor thin film [12], offering a unique and promising solution to the problems of material disorder and device scale-up.
Hyperdoping group IV semiconductors induces a superconducting state, yet routes to high quality, device-ready films have remained out of reach. Multiple studies have confirmed the existence of superconductivity [15, 2, 3, 16], however, these materials suffer from structural inhomogeneities, dopant clustering, and poor thickness control, which limit their applications. In terms of growth, the requirement of dopant concentrations above the standard miscibility limits raises concerns over the structural stability, site disorder, and nature of the superconducting phase [3]. Recent efforts in Ge have pushed dopant concentrations above miscibility limits, but maintaining coherent epitaxy and low defect densities remains challenging as most doping strategies rely on high-energy techniques such as ion implantation and flash annealing. These approaches demonstrate superconductivity but the formation of lattice defects and broad dopant profiles complicate reproducible film growth. This leads to persistent questions about whether these hyperdoped films are intrinsically superconducting, or if Ga clusters and unintended phases are responsible for superconductivity. These uncertainties highlight the need for growth strategies that introduce robust superconductivity into semiconductors without sacrificing crystallinity.
Figure 1: Superconductivity in germanium by p-type hyperdoping. A Schematic picture of dopants within a matrix and zone-center E(k) dispersion for a prototypical band insulator in a dilute doping and hyperdoped state. In the dilute doping scenario, dopants with an effective Bohr radius, $a_{0}$, at an average distance between dopants, $\lambda_{d}$, induce mid-gap electronic states at an energy $E_{a}$ above the valence band edge. Increasing dopant concentrations to few-percent levels or higher induce significant hybridization between dopants, giving rise to a low-energy band associated with dopant-dopant hybridization. This defect band is traditionally expected to be non-dispersive due to the stochastic behavior of dopants. However, dopant-dopant hybridization in principle can induce ordering of the dopants on the lattice, giving rise to a narrow heavy band as well. B and C present sheet resistance versus temperature traces for a single Ga:Ge layer and a Ga:Ge sandwich structure, akin to a Josephson junction along the growth direction, with a thin Si spacer layer, respectively. In both of these structures, the Ga:Ge hyperdoped layer is kept nominally constant, with the single layer thickness at roughly 12 nm and the Ga:Ge layer thicknesses in the sandwich are both 10 nm.
Here, we use molecular beam epitaxy (MBE) to grow Ga-hyperdoped Ge films, overcoming challenges associated with high dopant concentrations while maintaining coherent epitaxial growth. Crucially, the surfactant behavior of Ga enables the low-temperature growth of Ge, preserving smooth surfaces and single-crystallinity [17]. For films with an expected Ga doping concentration of 17.9%, we observe a superconducting transition at 3.5 K and a carrier concentration of 4.15 $\times$ 1021 holes cm-3. Employing short- and long-range structural characterization methods, we show that Ga substitution is energetically favored within the Ge lattice, slightly expanding its unit cell and producing a tetragonal distortion reminiscent of the $\beta$-Sn phase. Density functional theory (DFT) calculations reveal a pronounced shift of the Fermi level into the valence band and a flattening of electronic bands near the R-point of the cubic Brillouin zone. Our findings establish that superconductivity originates intrinsically in Ge from substitutional Ga doping and flat- or narrow-band physics, rather than defect-mediated superconducting weak links.
Figure 1A presents a schematic diagram of the dopant structure under dilute doping conditions, as well as a generalized picture for a hyperdoping in a semiconductor, along with their respective schematic E(k) dispersions. For dilute doping, the average distance between Ga dopants, $\lambda_{d}$, is larger than the effective Bohr radius of the Ga dopant, $a_{0}$. This results in discrete mid-gap states associated with these dopants at an energy $E_{\textup{a}}$ above the valence band edge, as is expected for a hole-doped semiconductor. However, the scenario in which dopant concentrations greatly exceed common doping levels, known as “hyperdoping”, leads to significant dopant-dopant and dopant-semiconductor hybridization of the low energy states. This gives rise to what is typically considered to be a non-dispersive band of defect states. Alternatively, hybridization with the host matrix can indeed give rise to a new heavy electronic band that disperses in $k$. Transport data for MBE-grown samples for this study is shown in Figure 1B and C, for structures containing a single layer of hyperdoped Ge and an epitaxial Josephson junction-like structure, respectively. Compared to previous reports on MBE-grown samples [16], these samples exhibit larger carrier concentrations that are more comparable to ion implantation strategies [15, 2, 3] (See Figure S1), albeit with a slight reduction in superconducting $T_{\textup{c}}$. We note a rough trend in $T_{\textup{c}}$ as a function of hole carriers where superconductivity appears near $1\times 10^{21}$ cm-3, reaching a maximum near $4-5\times 10^{21}$ cm-3. This is consistent with the idea that increasing the density of states (DOS) at the Fermi level increases the superconducting $T_{\textup{c}}$. Furthermore Figure 1B and C present the sheet resistance as a function of temperature for a 12 nm thick film of hyperdoped Ge and a sandwich heterostructure comprised of 10 nm layers of hyperdoped Ge surrounding a thin Si interlayer. The magnetic field dependence of longitudinal resistance within the Josephson structure is presented in Figure S3A and B for out-of-plane and in-plane field configurations, respectively. At 300 mK we observe the out-of-plane $H_{\textup{c}}$ to be 0.79 T, decreasing to 0.14 T at 2 K. In-plane $H_{\textup{c}}$ is found to be 2.51 T, decreasing to 0.78 T at 2 K. The robustness of this superconducting state in a magnetic field makes this superconducting system highly favorable for experiments investigating Andreev bound state behavior at values of magnetic field relevant for topological superconductivity [18, 19].
.1 Dopant fine structure and band structure
To examine the role of Ga as a dopant, we first study the thermodynamic preference and site-specific effects of introducing Ga into the diamond-type Ge crystal using first-principles calculations based on DFT, as detailed in the Methods section. Approximating from our Hall measurements and MBE growth conditions, doping concentrations of 12.5% and 11.1% are simulated for the case of substitutional and interstitial doping, respectively. Figure 2A shows the DFT-optimized structures. Incorporating group-III Ga as a heterovalent dopant within Ge implies some degree of non-equilibrium growth (i.e., the MBE growth process). This results in positive defect formation energies for cells containing Ga located at both doping sites (Figure 2A and Table S1). However, the formation energy of the substitutional Ga-doped structure (0.314 eV) is markedly lower than the interstitial positions (1.946 eV), indicating that substitutional dopants are expected to be more energetically favorable in the material. This preference is inherently connected to the requirement of the Ge host to undergo an expansion to accommodate the larger Ga atom at the interstitial site, as reported in Table S1. From our DFT calculations, substituting Ge with Ga at these doping levels is expected to impart a relatively small out-of-plane expansion of the lattice parameter of $\sim$0.12%, resulting in a slight tetragonal distortion to the structure (Table S1) to accommodate the heterovalent group-III dopant in a lattice-matched film.
Figure 2: Fine structure measurements revealing dopant locations and modulated band structure. A Optimized DFT structures of Ga-doped Ge at the interstitial and substitutional sites, along with the defect formation energies ($E_{\textup{Form}}$). B Experimental Ga-K-edge spectrum recorded from the Ga:Ge with its corresponding $k^{2}$-weighted $\chi$(k) EXAFS fit (red line) contained in the inset. C The Fourier-transform (k2-weighted) of the EXAFS data (distances have been phase-corrected). An agreeable fit has been made to the experiment using the substitutional doped crystallographic information file (CIF) generated from the optimized DFT calculated structure shown in A.D Calculated band structure for pristine Ge and the substitutionally doped GaGe7 structure. The horizontal dashed red line indicates the position of the Fermi level after doping, shifting 1.01 eV into the valence band. E The contribution of each band identified as i, ii, and iii to the 3D Fermi surface with the color map corresponding to the Fermi velocity. The predicted narrow-band condition is observed as small pockets of high-mass carriers in the surface of Band iii at the corners of the Brillouin Zone (R point).
To verify our theoretical calculations of doping site preference in hyperdoped Ge, we perform synchrotron X-ray absorption spectroscopy measurements. From our calculations (Figure S12), X-ray absorption at the Ga K-edge is inherently sensitive to the Ga-related local fine structure and can be used to unambiguously resolve its location within the lattice. Experimentally, we employ extended X-ray absorption fine structure (EXAFS) analysis, as measured from the Ga:Ge thin film in a fluorescent mode. Multiple spectra were recorded across the sample, with their average normalized spectrum forming the X-ray absorption data shown in Figure 2B. The $k^{2}$-weighted EXAFS signal is shown in the inset, with the corresponding analysis of the Fourier transform presented in Figure 2C. Considering the different structural environments that Ga can occupy in Ge (Figure 2A), the EXAFS data are best described by a substitutional doping model with a strong Ge-Ga nearest-neighbor scattering pair at 2.447 Å, which is close to the calculated bond length of 2.439 Å (Figure 2A). To rule out other possible coordination environments, such as $\alpha$-Ga segregation and the interstitial lattice position, Figure S12 contains the simulated profiles of these fine-structure environments. Together, with no strong fine structure observed between 2.5 and 3.8 Å in the experimental Fourier transform, these findings confirm that the Ga dopants preferentially occupy substitutional lattice positions, in agreement with our ab initio calculations. Furthermore, the relatively large Debye–Waller factors ($\sigma^{2}$( Å2)) derived from the EXAFS fits (Table LABEL:exafs-table) highlight the role of both lattice strain and Si diffusion in the Ga:Ge layers.
With confirmation of substitutional doping, we use the optimized structures in Figure 2A to study how the electronic structure is perturbed through relatively large incorporations of heterovalent Ga dopants at the different lattice sites. While a slight tetragonal distortion is anticipated in strained material due to substrate clamping, we find that the calculated band structure is insensitive to this distortion and we use the high-symmetry cubic Brillouin zone and paths for simplicity. The $k$-resolved band structure is presented in Figure 2D and the $k$-integrated DOS is in Figure S7A. From our formation energy calculations, Ga dopants are energetically unlikely to reside at interstitial lattice positions. This is confirmed by the DOS calculation shown in Figure S7A, which indicates that interstitial doping is undesirable for preserving the “clean” electronic bands of the host Ge semiconductor. This is evidenced by the disappearance of the parent Ge band gap and the emergence of metallic states. In contrast, substitutional Ga doping is expected to form acceptor states below the valence band edge, while the overall band structure near the gap remains largely unperturbed (Figure 2D). An evaluation of the DOS projected onto different atomic orbitals in Figure S7B shows that new, localized states emerge through the hybridization of both the Ga and Ge p-orbitals. As expected, substituting Ge for group III Ga induces p-type doping in the system, with the Fermi level shifting 1.01 eV into the valence band (see Figure 2D). This corresponds to a hole carrier concentration of $5.5\times 10^{21}$ cm-3 that closely approximates the hole concentration from our Hall measurements. Furthermore, due to the lifting and flattening of the bands near the R point of the cubic Brillouin zone, we predict this system to be a candidate narrow-band superconductor. By plotting the Fermi surface and separating the contribution of each of the three bands crossing the Fermi level (band i, ii, and iii), we observe in the surface for band iii, hole pockets of high-mass carriers that cross the Fermi level at the corners of the Brillouin zone (Figure 2E). This feature is clearly represented in the 2D projected cross-sections of the doped Ga:Ge Fermi surface (Figure S8) and is highly sensitive to the position of the Fermi level. Lowering it by 40 meV to simulate a slight increase in the hole concentration will both enlarge the corner hole pocket projections for band iii and introduce similar features into the projected corners of band ii (Figure S9).
Figure 3: Cross-sectional electron microscopy reveals coherent crystalline interfaces. A Schematic of the film structure as grown by MBE. B-C Cross-sectional TEM/EDS imaging of the $T_{\textup{c}}$ = 3.5 K hyperdoped sample presenting the compositional profile across the entire film thickness. D Low-magnification cross-sectional HAADF-STEM image of the film. E Zoom-in on the film/substrate and Ga:Ge/Si interfaces, displaying coherent crystalline interfaces. F Raman back-scattering spectra of undoped (black) and doped (red) structures using 532 nm laser excitation at room temperature. The inset highlights the common emergence of highly anharmonic second-order optical bands, identified as the active LO mode here due to the Raman selection rules for backscattering from a (001) Ge surface.
.2 Epitaxial interfaces and Phonons
Figure 3A presents a schematic structure of the grown film with $T_{\textup{c}}$ = 3.5 K, in which two hyperdoped Ge layers are grown sandwiched between a thin Si buffer layer. Cross-sectional transmission electron microscopy (TEM) and energy dispersive spectroscopy (EDS) measurements are presented in Figure 3B-C and show the spatially homogeneous elemental profile of the Ga dopants throughout the hyperdoped layer. Notably, no distinct signatures of Ga clustering are observed compared to earlier reports where annealing was used to activate dopants  [3]. Likewise, these MBE-grown materials also exhibit abrupt interfaces and a smooth surface amenable to future planar device fabrication as opposed to previous MBE studies [16]. A view of the atomic structure using low-magnification high-angle annular dark field scanning transmission electron microscopy (HAADF-STEM) is presented in Figure 3D, with a corresponding expansion of the key film/substrate and Ga:Ge/Si interfaces contained in Figure 3E. Owing to the small amount of mismatch strain and thin layer thicknesses, the growth of Ga:Ge is pseudomorphic (commensurate) with the underlying Ge substrate and the atomic planes on both sides of the interface are in perfect registry. This demonstration of fine thickness control, commensurate interfaces, and homogeneous dopant distribution exhibits all of the necessary structural requirements for the fabrication of low-disorder superconducting devices that have been proposed to mitigate loss channels in quantum information applications [12].
Evaluating the (zone center) phonon spectra of pure Ge and hyperdoped Ge films using Raman backscattering spectroscopy (Figure 3F), the high p-type concentration produces a new, softened mode near 280 cm-1 [20]. A consequence of this mode is the screening of the Ge-Ge optical mode, which is now coupled to the large hole concentration and explains its significant reduction in the doped specimen. Lattice dynamics calculations (Figure S11) of pure and hyperdoped systems support this observation and reveal phonon softening in hyperdoped Ge. Evaluating the Raman spectra in Figure 3F, both doped and undoped samples yield comparable Raman signals related to the Si-Si and Si-Ge-type sublattices. For the Si-Si optical vibration originating from the thin Si layers, the low dimensionality and diffusion of Ge and Ga atoms into this region lead to a significant mode softening and a strong Si-Ge band near 400 cm-1 (Figure 3C) [21]. As both doped and control films exhibit similar Si-Si, Si-Ge, and Ge-Ge bands, we infer high Ga doping does not appreciably affect the growth process.
Figure 4: A 2D GISAXS pattern of the Ga:Ge epitaxial film recorded at the critical angle (0.26) of the superconducting layer. The intensity map is on a linear scale. B The out-of-plane X-ray reflectivity measurement of the Ga:Ge film with a fit made using the REFLEX software package [22]. C Cross-section of the electron scattering length density (ESLD) of the epitaxial stack derived for the fitting model of the fringes in B. D 2D GIWAXS pattern of the Ga:Ge thin film record at the critical angle (0.125). The color map is on a log scale. The family of crystallographic planes giving rise to the intense Bragg reflections of the hyperdoped layer are identified with arrows. E Corresponding integrated GIWAXS profile (qxyz) and its subsequent structural refinement (Le Bail method). The refined unit cell parameters are displayed and an expansion of the split 220 family of Bragg peaks is shown in the inset to highlight the reduced symmetry of the clamped Ga:Ge epitaxial layer in comparison to the undoped Ge control.
.3 Observation of crystalline distortion
We next resolve the microstructure of the hyperdoped heterostructure using a combination of synchrotron-based grazing incidence X-ray scattering experiments. The 2D grazing-incidence small-angle x-ray scattering (GISAXS) pattern presented in Figure 4A was recorded at the critical angle of the Ga:Ge layers (0.26 ), which is well above the critical angle of the relatively less dense Si layers (Figure S12). Pronounced Keising fringes directed in the out-of-plane direction emerge due to the interference of scattered photons within the Ga:Ge epitaxial film and are evidence of sharp interfaces between the layers. We employ X-ray reflectometry (XRR) to derive the electron density profile and compositional grading parameters (rms roughness) of the film in the normal direction. Based on the XRR fit shown in Figure 4B, compositional grading is confirmed between the layers and diffusion is revealed to be asymmetrical with respect to the stack, with larger roughness values determined for Si grown on top of Ga:Ge, compared to Ga:Ge grown on the Si layer.
[…]
.4 Conclusions
In conclusion, our results establish that MBE growth of hyperdoped Ga:Ge exhibits superconductivity arising from substitutional Ga incorporation, allowing for the creation of coherent semiconductor-superconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal ($\beta$-Sn) distortion, in line with our DFT results. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level deep into the valence band and flattens the electronic bands near the R point, likely promoting superconductivity. Our findings highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge and underscore its potential as a scalable platform for epitaxial semiconductor-superconductor heterostructures using the mature group IV semiconductor toolkit.
Acknowledgments
S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation. P.J. acknowledges fruitful discussions with Ross McKenzie and Ben Powell.
Funding:
P.J.S and J.S. acknowledge funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the NCI National Facility through the National Computational Merit Allocation Scheme. This work was partially supported by the Australian Research Council under the following grants: DE230100173 (J.A.S.), LP210200636 (Y.-H.C., P.J.), DE220101147 (C.V.). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University.
Author contributions:
Conceptualization: P.J., J.S., and E.D. Methodology: J.A.S., P.J.S., C.V., S.S.-R. Investigation: J.A.S., P.J.S., C.V., A.B., A.D., Y.-H.C., J.V.D., S.S.-R. Visualization: J.A.S., P.J.S. Funding acquisition: J.A.S., P.J., J.S. Project administration: J.A.S., P.J.S., P.J., J.S. Supervision: J.A.S., P.J., J.S., E.D., L.W. Writing – original draft: P.J., J.A.S., P.J.S. Writing – review and editing: J.A.S., P.J.S., C.V., P.J., J.S.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Raw data available on Zenodo: doi.org/10.5281/zenodo.14782527
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S14
Tables S1 to S2
References (23-0)
Supplementary Materials for
Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics
Julian A. Steele∗†, Patrick J. Strohbeen, Carla Verdi, Ardeshir Baktash,
Alisa Danilenko, Yi-Hsun Chen, Jechiel van Dijk, Lianzhou Wang,
Eugene Demler, Salva Salmani-Rezaie, Peter Jacobson, Javad Shabani
Corresponding author. Email: julian.steele@uq.edu.au, jshabani@nyu.edu
These authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S14
Tables S1 to S2
Materials and Methods
Thin film growth: The materials used in this study are all grown in a custom Varian Gen II molecular beam epitaxy (MBE) system on 50.8 mm undoped Ge (001) wafers. Prior to loading into vacuum, the wafers are etched in deionized water at 90 C for 15 minutes and then immediately loaded into the vacuum chamber on indium-free mounting blocks. The wafers are sequentially outgassed inside the growth reactor at 250 C, 450 C, and finally 650 C, for 15 minutes, 15 minutes, and 5 minutes, respectively, then cooled to room temperature for growth. Pure germanium (6N) and silicon (3N) source material is deposited via Thermionics HM2 e-Guns operating at a 6 kV acceleration voltage. Gallium doping is done with a standard Knudsen effusion cell source (MBE Komponenten). Atomic fluxes are measured prior to growth with a retractable quartz crystal microbalance placed in the center of the growth path in front of the substrate. The substrate is faced away from the atomic beam path(s) during flux measurement. Reported temperatures are from a thermocouple attached close to the backside of the substrate that has been calibrated to the GaAs (001) oxide-removal temperature. Samples are grown with a Ge flux of $\sim 8-8.5\times 10^{13}~{}cm^{-2},s^{-1}$, a Ga flux of $1.4-1.9\times 10^{13}~{}cm^{-2},s^{-1}$, and a Si flux of $2.5-3\times 10^{13}~{}cm^{-2},s^{-1}$. The nominally 20 nm thick films are grown as follows, as described in Ref. [16] : (1) 10 nm of Ga-doped Ge, (2) 0.5 nm of Si, (3) 10 nm of Ga-doped Ge, (4) 1 nm Si cap. Film growth is monitored using reflection high-energy electron diffraction (RHEED) along a $\langle 110\rangle$ cut (beam path $\parallel$ to $\langle 010\rangle$). Representative RHEED images throughout the growth steps are presented in Figure S1.
Four-point transport measurements: Electrical characterization is done in an Oxford TeslatronPT He4 cryostat using a HelioxVT He3 probe insert with a base temperature of $\sim$ 270 mK and magnetic field capabilities up to 12 T. Measurements are collected using a standard Van der Pauw wiring configuration on square pieces cleaved from near the center of each wafer. On-chip contacts are made via annealed In-Sn eutectic alloy at each of the four corners which are then wired to a daughterboard using gold wire.
First-principles calculations: First-principles calculations were performed using DFT as implemented in VASP [25, 26]. Projector augmented wave (PAW) pseudopotentials including semicore 3d electrons were adopted, with a plane-wave cutoff of 550 eV. A $7\times 7\times 7$ $\Gamma$-centered k-point mesh was used to sample the Brillouin zone of the conventional 8-atom unit cell of Ge (cubic diamond structure, space group $Fd\overline{3}m$). Structural optimizations were performed using the PBEsol exchange-correlation functional [27], which yields lattice parameters in excellent agreement with the experiment (see Table S1), and a convergence threshold of 0.005 eV/Å on all forces.
For the Ga-doped Ge structure with substitutional fine-structure, one host atom in the 8-atom unit cell is replaced by a Ga atom, corresponding to a doping concentration of 12.5%. Conversely, interstitial doping is modeled by placing one Ga atom at interstitial doping sites, representing a doping concentration of 11.1%. To model the effect of substrate clamping in Ga-doped Ge, the in-plane lattice parameters were fixed to the pristine Ge values, and only the out-of-plane lattice parameter was allowed to relax. K-edge spectra were simulated using the supercell core-hole method [28] within a $2\times 2\times 2$ supercell containing 64 atoms (72 atoms in the case of interstitial doping). The formation energy of a neutral Ga dopant in the Ge crystal is calculated as $E_{\textup{Form}}=E_{\textup{d}}-E_{\textup{p}}-\sum_{i}n_{i}\mu_{i}$, where $E_{\textup{d}}$ is the total energy of the structure containing the defect, $E_{\textup{p}}$ is the total energy of the pristine structure, $n_{i}$ is the number of atoms of species $i$ being added ($n>0$) or removed ($n<0$), and $\mu_{i}$ is the chemical potential. For Ga, this is the energy per atom of a Ga crystal (base-centered orthorhombic structure).
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional [29] based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge [30, 31], which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included in all electronic structure calculations. A denser $20\times 20\times 20$ k-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a $31\times 31\times 31$ k-mesh was used. Vibrational properties were computed using finite differences [32] within supercells containing 216 atoms ($3\times 3\times 3$ supercells of the 8-atom unit cell), employing $3\times 3\times 3$ $k$-point meshes and displacement amplitudes of 0.015 Å.
Scanning transmission electron microscopy: Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64-200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 $\times$ 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy: Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal micro-Raman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10x, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 cm-1), instrument calibration was verified through checking the position of the Si band at $\pm$520.7 cm-1, and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S10). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
Synchrotron-based GISAXS/GIWAXS: To identify the crystalline phases and microstructure of films, synchrotron-based grazing incidence wide-angle X-ray scattering (GIWAXS) data were collected at the small angle X-ray scattering/wide angle X-ray scattering (SAXS/WAXS) beamline at the Australian Synchrotron [33]: 2D scattering patterns were recorded using a wavelength of either 1.340375 Å (9.25 keV) or 0.619924 Å (20 keV), using an energy-selective Pilatus 2M detector. The measurements made using a 20 keV beam energy were to increase the observed Q-range for the structural refinements, using an energy threshold of 15 keV to suppress the X-ray fluorescence signals arising from the excited K-edges of Ge (11.1 keV) and Ga (10.37 keV). Experiments at 9.25 keV had no such issues. Two different sample-to-detector distances were employed to record both SAXS (5.5 m) and WAXS (0.45 m) profiles, each calibrated using a silver behenate reference standard. The sample and detector were enclosed in a vacuum chamber to suppress air scatter. Scattering patterns were measured as a function of the angle of incidence, with data shown acquired with an angle of incidence near the critical angle (for a given X-ray energy) to maximize scattering intensity from the sample. All collected 2D images were azimuthally integrated using PyFAI [34] and processed using a custom Python routine. The resulting unit cell models were refined using the La Bail method implemented in Fullprof [35], a comprehensive analysis software. In line with our other characterization techniques, our GIWAXS experiments failed to detect Ga droplet formation or segregation within the film. Due to the relatively small scattering volume of the Si layers, no additional peaks are indexed to this phase either.
Stress analysis: The lattice mismatch is accommodated by an elastic strain in the hyperdoped epilayer, giving a biaxial stress of $\sigma=2\mu f\cdot(1+\nu)/(1-\nu)$, where $\mu$ is the shear modulus, $\nu$ Poisson’s ratio and $f$ the misfit parameter. Assuming the hyperdoped Ga:Ge epilayer forms a cubic unit cell when unconstrained, elastic isotropy is assumed in the relevant deformation directions and the misfit parameter is given by $f=(a_{Ga:Ge}-a_{Ge})/a_{Ge}$, where $a_{Ga:Ge}$ and $a_{Ge}$ are the lattice parameters of the relaxed epilayer and substrate, respectively. We do not have such experimental data for the relaxed hyperdoped unit cell, however, since the epilayers studied here are pseudomorphically grown (Figure 3e), the values obtained from the structural refinement can be related to the relaxed lattice parameter of the Ga:Ge layer through $a_{Ga:Ge}=a_{Ge}[1+P(\Delta d/d)\perp$], where P is the elastic parameter of the films [36]. Regardless of the growth direction, the in-plane and out-of-plane biaxial strain components are given by $\epsilon_{\parallel}=(a_{Ga:Ge}-a_{Ge})/a_{Ge}$ and $\epsilon_{\perp}=\epsilon_{\parallel}(1-1/P)$, respectively. For our thin films grown on Ge (100); $P_{100}=C_{11}/(C_{11}+2C_{12})=0.5709$. To implement our calculation, we approximate the elastic constants $C_{ij}$ of the hyperdoped film to be comparable to the pure Ge crystal [36] (as there is no relevant data for GaGe alloys). Based on our hole carrier concentrations, the Ga concentration for the sample studied in Figure 3e is $\sim$11.22and has a calculated relaxed (cubic) lattice parameter of $a_{Ga:Ge}$ = 5.6614 Å, which is 0.11% larger than the Ge crystal.
Synchrotron-based EXAFS experiments: Ga K-edge X-ray absorption spectra (XAS) were captured at the Australian Synchrotron (ANSTO in Clayton, Victoria) MEX1 beamline. A double-crystal Si (111) monochromator equipped with focusing optics was used to reduce harmonic content while producing excitation energy. An inline Ga2O3 reference (the initial peak of the first derivative occurs at 10.3687 keV) was used to calibrate the monochromator at Ga-K absorption edge with E0 set to its known value of 10.3671 eV. All acquisitions were performed on thin film materials in a fluorescent mode. Brief XAS scans of specific samples were conducted to verify the stability of the materials under at least several minutes of X-ray exposure before the full acquisitions. Samples were measured with varying energy intervals over the pre-edge (5 eV) and the XANES region (0.25), with 0.035 intervals in k-space over the EXAFS. At low-k an integration time of 1 s was employed per interval, with longer integration times weighted toward high-k portions of the spectrum, up to a maximum value of 9 in k-space (max of 2 s). Multiple scans were collected at different positions of the sample. Data processing including background subtraction, scan averaging, edge-height normalization, and rebinning was performed based on the Athena program [37]. For normalization of energy spectra and removal of background, the pre-edge range was set between -150 eV and -30 eV, while the normalization range spanned from 80 to 300 eV post-edge. The order of normalization was designated as 3. For the presented EXAFS data, the k-weight was configured as 2, and the k-range for the forward Fourier transform was defined from 0.8 to 8.5.
EXAFS fine-structure modeling: Fits was conducted in the ARTEMIS, part of the IFEFFIT software package [37]. The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package [37], with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table LABEL:exafs-table). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.
[…]
Figure S1: Comparison of this study against literature reports for ion implantation superconducting TC as a function of Hall carriers. The colored star, square, and triangular points were prepared via ion implantation [15, 2, 3] while the black circles (this study) are MBE-grown samples
Figure S2: RHEED line cuts for superconducting hyperdoped Ge thin film. A Ge substrate $\langle 110\rangle$-type direction, taken at room temperature just before deposition of Ga:Ge film. B Initial surface of Ga:Ge film. C RHEED pattern after the first 10nm of Ga:Ge deposited. D After the Si spacer layer and initial growth of the second Ga:Ge layer. E End of Ga:Ge growth, before Si cap is deposited. We note a superimposed amorphous ring over a faint $\langle 110\rangle$-type direction diffraction pattern. This is consistent with previous observations that Ga-adsorption at low temperatures creates a floating layer at the surface from which the incorporated dopants are taken [17].
Figure S3: Out-of-plane A and in-plane B critical magnetic field behavior for the Ga:Ge sandwich heterostructure presented in Figure 1C and Figure 3.
Figure S4: Secondary-Ion Mass Specroscopy data for two samples investigated in this study, A PJS146 ($T_{\textup{c}}\sim$ 2.5 K) and B PJS147 ($T_{\textup{c}}\sim$ 3.5 K))
Figure S5: Calculated X-ray absorption spectra of Ga K-edge for Ga:Ge crystals with substitutional and interstitial doping sites.
Figure S6: Comparison of the simulated fine structure derived from the CIF files of Ga:Ge interstitial system (DFT calculated system in Figure 2A of main text) and amorphous Ga metal [38] Note that these Fourier transforms have been phase corrected, for simplicity. The arrows highlight the relatively large difference in experimental bond length of the nearest-neighbor scattering pair and those derived in the simulations.
Figure S7: A Total density of states (DOS) calculated for the pure Ge crystal and the hyperdoped Ga:Ge shown in Fig. 2A of the main text, with Ga at interstitial (int.) and substitutional (sub.) lattice positions. The energies are aligned to the valence band maximum (VBM), set as the zero of the energy. The Fermi level in each case is indicated by a dashed vertical line. B Total and projected DOS of the hyperdoped Ga:Ge crystal with substitutional fine-structure.
Figure S8: A 3D Fermi surface for substitutionally doped Ga:Ge, showing the (011) plane. B 2D Fermi surface cut on the (011) plane, showing the contribution of bands i, ii, and iii. Hole pockets from band iii are visible in the R-point corners.
Figure S9: Fermi surface obtained after shifting the Fermi level by 40 meV below the value shown in Figure 2B of the main text. A Contribution of bands i, ii, and iii to the 3D Fermi surface, with the color corresponding to the Fermi velocity. B (011) plane crossing the 3D Fermi surface. C 2D cut on this plane, showing the contribution of each band. The predicted narrow- or flat-band conditions are shown to be sensitive to the position of the Fermi level, as the pockets of high-mass carriers from bands ii and iii at the corners of the Brillouin Zone (R point) are shown to enlarge with a relatively minor shift in the Fermi level.
Figure S10: A Raman Scattering penetration depth ($\mathrm{d_{\mathrm{RS}}}$) as a function of the laser excitation energy, determined by considering the absorption coefficients [39] for both the incoming ($\alpha_{\mathrm{out}}$) and the outgoing ($\alpha_{\mathrm{in}}$) light: $\mathrm{d_{\mathrm{RS}}}=1/(\alpha_{\mathrm{in}}+\alpha_{\mathrm{out}})\approx 1/\alpha_{in}$. B Comparison of Raman spectra recorded from Ga:Ge and Ge control MBE films at different excitation energies. At longer excitation wavelengths (785 nm), both the control and target films exhibit a comparable Raman response. Probing the surface-sensitive information with shorter wavelengths, the differences between the vibrational structure of the films become clearer. The appearance of disorder-activated LA modes below 200 cm-1 (from both Ge-type and Si-type sublattices) in the scattering volume yields comparable scattering intensities under these conditions for the control and doped sample. We suggest the origins of these bands are a combination of (local) lattice strain, phonon confinement and (partial) satisfaction of the resonance conditions of the Ge crystal $E_{1}+\Delta_{1}$ gap [40]. Together, these features break the normal Raman scattering selection rules of a normal (001)-oriented bulk Ge film, i.e., resembling more the Raman spectra recorded using 785 nm laser light.
Figure S11: Calculated first-order phonon DOS for pure Ge and hyperdoped Ga:Ge crystal.
Figure S12: The calculated penetration depth of the different material layers at the synchrotron energies used for grazing incidence experiments. For the range of angles below and at the critical angle, the X-ray electromagnetic field only interacts a short distance below the film surface due to the evanescent damping (5 - 10 nm), and constitutes the defined evanescent regime.
Figure S13: Evolution of the integrated GIWAXS signal recorded from the Ga:Ge film as a function of the grazing incidence angle ($\alpha$i: values inset). Here * identifies the calculated critical angle for the doped Ga:Ge layer at an X-ray beam energy of 20 keV. As the angle of incidence nears the critical angle of the doped layer, the Bragg reflections emerging begin to intensify. A combination of lateral steps (200 $\mu$m) made between each incident angle frame, mechanical jolting, and imperfect beam divergence means the critical angle conditions (waveguide type effect) is relatively really sensitive and results in the discontinuity of some profile features between frames.
Figure S14: Integrated GIWAXS profile (qxyz) and structural refinement (Le Bail method) of the pure Ge epitaxial layer. Unstrained Ge crystal belongs to the high-symmetry cubic space group $Fd\overline{3}m$ (space group No. 227), characterized by equivalent lattice parameters (a=b=c) and a structural refinement of our undoped Si-Ge film aligns well with the reported bulk, cubic value (a = 5.657906$\pm$0.0000009 Å, Vol. = 181.1203 Å3)[41]
Table S1: Structural parameters of DFT optimized structures and defect formation energies $E_{\textup{Form}}$. Values are calculated for pure and Ga-doped Ge (both substitutional and interstitial doping).
Structure Lattice Parameter (Å) Volume (Å3) Crystal Structure $E_{\textup{Form}}$ (eV)
a=b, c
Pure Ge 5.674, 5.674 182.64 Cubic
Sub. Ga:Ge 5.674, 5.681 182.87 Tetragonal 0.314
Int. Ga:Ge 5.674, 6.220 200.128 Tetragonal 1.946
Table S2: EXAFS fitting parameters. Values are extracted from fits made to spectra recorded at 295 K using a Ga substitutional model. The lower ranked paths (in italics) are omitted from the actual fit model to avoid overfitting. The errors provided in the parentheses align with the last significant figures of the variable fit values. Further fitting details can be found in the Methods section. The path type can be tracked as: SS - single scattering pair; DS - double scattering; R - rattle.
Path type Degen. Rank R(Å $\sigma^{3}$2) $\Delta$E0 (eV) S02 R-factor
Ga-Ge (SS) 4 100 2.451(1) 0.0051(4)
Ga-Ge (SS) 12 90.82 4.002(5) 0.0063(2)
Ga-Ge-Ge (DS) 12 5.71 4.450
Ga-Ge-Ge (DS) 24 21.4 4.452(2) 0.010(6) -0.195 0.68(11) 0.015
Ga-Ge (SS) 12 59.45 4.693(2) 0.0087(5)
Ga-Ge-Ge (R) 4 4.09 4.900