### v3

# Superconductivity in Substitutional Ga-Hyperdoped Ge Epitaxial Thin Films

Abstract
Doping-induced superconductivity in group IV elements may enable quantum functionalities in material systems accessible with well-established semiconductor technologies. Non-equilibrium hyperdoping of group III atoms into C, Si, or Ge can yield superconductivity; however, its origin is obscured by structural disorder and dopant clustering. Here, we report the epitaxial growth of hyperdoped Ga:Ge films and trilayer heterostructures by molecular beam epitaxy with extreme hole concentrations ($n_{\textup{h}}=4.15\times 10^{21}$ cm-3,  17.9% Ga substitution) that yield superconductivity with a critical temperature of $T_{\textup{c}}=3.5$ K. Synchrotron-based X-ray absorption and scattering methods reveal that Ga dopants are substitutionally incorporated within the Ge lattice, introducing a tetragonal distortion to the crystal unit cell. Our findings, corroborated by first-principles calculations, suggest that the structural order of Ga dopants creates a narrow band for the emergence of superconductivity in Ge, establishing hyperdoped Ga:Ge as a low-disorder, epitaxial superconductor-semiconductor platform.
Realization of superconductivity in group IV elements has been an enticing field of study over the past few decades for their promising application in superconducting electronics Ekimov et al. (2004); Bustarret et al. (2006); Blase et al. (2009); Chiodi et al. (2017) such as superconducting quantum bits or cryogenic CMOS control circuitry. However, difficulties such as dopant segregation and precipitation, nanocrystal formation, incoherent interfaces, and poor layer thickness control impede their integration into a new generation of quantum devices. As a prototypical semiconductor and foundry-ready material, germanium is a key component of devices that combine macroscopic superconducting coherence with the microscopic degrees of freedom in semiconductors Scappucci et al. (2021); Sammak et al. (2019); Tosato et al. (2023); Schiela et al. (2024). Historically, superconductivity in germanium has been accomplished through extreme non-equilibrium growth techniques such as ion implantation and subsequent flash annealing Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021). However, recent advancements in thin film epitaxy of heavily doped germanium thin films have shown great promise in achieving well-ordered, single-crystalline thin films of superconducting germanium Strohbeen et al. (2023). Homoepitaxy of superconducting group IV materials allows for expanding upon the existing repertoire of superconductor-semiconductor (S-Sm) planar devices Shabani et al. (2016); Krogstrup et al. (2015) to enable full S-Sm-S heterostructures that have remained challenging to realize, thus limiting their applicability in quantum devices Zhao et al. (2020); Shim and Tahan (2014). This difficulty has renewed interest in doping-based superconductivity in group IV materials, a concept explored theoretically by Cohen in the 1960s Cohen (1964), as recent efforts to induce superconductivity through proximity effects Tosato et al. (2023); Schiela et al. (2024) and alloying Strohbeen et al. (2024) have met with mixed results. This route sidesteps disordered interfaces between dissimilar materials by inducing superconductivity within an epitaxial semiconductor thin film Shim and Tahan (2014), offering a unique and promising solution to the problems of material disorder and device scale-up.
Hyperdoping in germanium induces a superconducting state, yet routes to high quality, device-ready films have remained out of reach. Multiple studies have confirmed the existence of superconductivity Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021); Strohbeen et al. (2023), however, these materials suffer from structural inhomogeneities, dopant clustering, and poor thickness control, which limit their applications. Specifically, doping at concentrations above the standard miscibility limits raises concerns over the structural stability, site disorder, and nature of the superconducting phase Sardashti et al. (2021). Recent efforts in Ge have pushed dopant concentrations above miscibility limits Strohbeen et al. (2023); Trumbore (1960); Sardashti et al. (2021); Prucnal et al. (2019); Hermannsdörfer et al. (2009), but maintaining coherent epitaxy and low defect densities remains challenging as most doping strategies rely on high-energy techniques such as ion implantation and flash annealing. These approaches demonstrate superconductivity but the formation of lattice defects and broad dopant profiles complicate device integration and dopant uniformity. Furthermore, questions persist about whether these hyperdoped films are intrinsically superconducting, or if Ga clusters and unintended phases are responsible for superconductivity. These uncertainties highlight the need for growth strategies more readily amenable to achieve the quality necessary for quantum devices as well as better understanding of dopant incorporation in these extreme doping regimes.
Here, we use molecular beam epitaxy (MBE) to grow Ga-hyperdoped Ge films, overcoming challenges associated with high dopant concentrations while maintaining coherent epitaxial growth. Crucially, better control over parasitic heating during film growth enables the formation of a smooth floating layer of Ga adatoms that promotes low-temperature growth of Ge, preserving smooth surfaces and single-crystallinity Kesan et al. (1991). By re-configuring the growth chamber and optimizing growth temperatures, we obtain drastic improvements in surface morphology, atomic ordering, and domain structure. We demonstrate the improved material quality in the growth of trilayer heterostructures as a proof-of-principle design for vertical Josephson junction device architectures that alleviate concerns of material disorder related to typical amorphous oxide tunnel junctions. For films with an expected Ga doping concentration of 17.9%, we observe a superconducting transition at 3.5 K and a carrier concentration of $4.15\times 10^{21}$ holes cm-3. Short- and long-range structural characterization methods show that Ga substitution is energetically favored within the Ge lattice and drives a tetragonal distortion reminiscent of the $\beta$-Sn phase. Density functional theory (DFT) calculations reveal a pronounced shift of the Fermi level into the valence band and a flattening of electronic bands near the R-point of the cubic Brillouin zone. Our findings establish that superconductivity originates intrinsically in Ge from substitutional Ga doping and flat- or narrow-band physics, rather than defect-mediated superconducting weak links.
Figure 1: Superconductivity in germanium by p-type hyperdoping. a Schematic picture of dopants within a matrix and zone-center $E(k)$ dispersion for a prototypical band insulator in a dilute doping and hyperdoped state. In the dilute doping scenario, dopants with an effective Bohr radius, $a_{0}$, at an average distance between dopants, $\lambda_{d}$, induce mid-gap electronic states at an energy $E_{a}$ above the valence band edge. Increasing dopant concentrations to few-percent levels or higher induces significant hybridization between dopants, giving rise to a low-energy band associated with dopant-dopant hybridization. This defect band is traditionally expected to be non-dispersive due to the stochastic behavior of dopants. However, dopant-dopant hybridization in principle can induce ordering of the dopants on the lattice, giving rise to a narrow heavy band as well. b-e Sheet resistance versus temperature traces for single Ga:Ge layers and Ga:Ge sandwich structures, akin to a Josephson junction along the growth direction. In these structures, the single Ga:Ge layers are held to a constant thickness of roughly 12 nm and the trilayer structures have Ga:Ge layer thicknesses of 10 nm. The insets of each panel present schematic drawings of the sample layer structure. b, c are representative traces where we use Si as the cap and barrier layer material, and d, e are representative traces where Ge replaces the Si-containing layers. The jitter in temperature observed in d during the superconducting transition is an artifact caused by temperature instability during the measurement.
Thin film growth and material stability
Figure 1a presents a schematic diagram of the dopant structure under dilute doping conditions, as well as a generalized picture for hyperdoping in a semiconductor, along with their respective schematic $E(k)$ dispersions. For dilute doping, the average distance between Ga dopants, $\lambda_{d}$, is larger than the effective Bohr radius of the Ga dopant, $a_{0}$. This results in discrete mid-gap states associated with these dopants at an energy $E_{\textup{a}}$ above the valence band edge, as is expected for a hole-doped semiconductor. However, the scenario in which dopant concentrations greatly exceed common doping levels, known as “hyperdoping”, leads to significant dopant-dopant and dopant-semiconductor hybridization of the low-energy states. This gives rise to what is typically considered to be a non-dispersive band of defect states. Alternatively, hybridization with the host matrix can indeed give rise to a new heavy electronic band that disperses in $k$. Transport data for MBE-grown samples for this study are shown in Figure 1b and c, for structures containing a single layer of hyperdoped Ge and an epitaxial Josephson junction-like structure, respectively. Compared to previous reports on MBE-grown samples Strohbeen et al. (2023), these films exhibit a slight reduction in superconducting $T_{\textup{c}}$, but larger carrier concentrations that are more comparable to ion implantation strategies Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021) (see Figure S1a). Figure S1a shows a rough trend in $T_{\textup{c}}$ as a function of hole carriers where superconductivity appears near $1\times 10^{21}$ cm-3, reaching a maximum near $4-5\times 10^{21}$ cm-3 in the samples grown for this study. This is consistent with a previous report in boron-doped silicon Nath et al. (2024) that demonstrates an increasing hole density increases the superconducting $T_{\textup{c}}$. Furthermore, Figure 1b and c present the sheet resistance as a function of temperature for a 12 nm thick film of hyperdoped Ge and a trilayer heterostructure, respectively. Representative reflection high-energy electron diffraction (RHEED) images for a layered heterostructure of hyperdoped Ge films are presented in Figure S2. During Ga doping, we observe the formation of the expected floating surface layer of Ga adatoms Kesan et al. (1991) with an underlying single crystalline pattern that reflects the surface structure of the hyperdoped film. Further details regarding RHEED monitoring and film growth are presented in the Supplementary information and Methods, respectively. The smoothness of the resultant film surface is confirmed with ex-situ atomic force microscopy (AFM) imaging, seen in Figure S1b.
To investigate the role of the silicon interlayer more explicitly, films grown without silicon were also studied. The results of which are presented in Figures 1d, e, and S3. We note that the silicon-containing samples stabilize slightly higher carrier concentrations, which we expect to exist due to the band offset between the hyperdoped Ge and strained Si layers Mönch (1987); Van de Walle (1988). Silicon incorporation also enhances the resultant superconducting $T_{\textup{c}}$, which we attribute to a behavior akin to the BCS isotope effect Blase et al. (2009); Nath et al. (2024). Temperature stability of the superconducting phase is shown in Figure S3, where we obtain a maximal superconducting transition temperature at growth temperatures ranging from 100-150C. Above 150C the superconducting phase weakens as observed by a reduction in $T_{\textup{c}}$, coinciding with a reduction in hole carriers seen in Figure S4. We further investigate the robustness of the superconducting phase by fabricating a Hall bar device out of the sample shown in Figure 1c. We present the data for this test device in Figure S5 and the fabrication procedure is found in the Methods. We note no significant change in superconducting $T_{\textup{c}}$, out-of-plane $H_{\textup{c}}$, nor carrier concentration. An extended discussion of this data is provided in the Supplementary Information. Thus, we conclude that this superconducting phase is demonstrably stable during processing and suitable for device applications.
Figure 2: Ga dopant arrangement in Ge and modulated band structure. a Optimized DFT structures of Ga-doped Ge at the interstitial and substitutional sites, along with the defect formation energies ($E_{\textup{Form}}$). b Experimental Ga-K-edge spectrum recorded from the Ga:Ge with its corresponding $k^{2}$-weighted $\chi(k)$ EXAFS fit (red line) contained in the inset. c The Fourier-transform ($k^{2}$-weighted) of the EXAFS data (distances have been phase-corrected). An agreeable fit has been made to the experiment using the substitutional doped crystallographic information file (CIF) generated from the optimized DFT calculated structure shown in a.d Calculated band structure for pristine Ge and the substitutionally doped GaGe7 structure. The horizontal dashed red line indicates the position of the Fermi level after doping, shifting 1.01 eV into the valence band. e The contribution of each band identified as i, ii, and iii to the 3D Fermi surface with the color map corresponding to the Fermi velocity. The predicted narrow-band condition is observed as small pockets of high-mass carriers in the surface of Band iii at the corners of the Brillouin Zone (R point).
Dopant arrangement and band structure
Approximating from our Hall measurements and MBE growth conditions, doping concentrations of 12.5% and 11.1% are simulated for the case of substitutional and interstitial doping, respectively, as detailed in the Methods section. Figure 2a shows the DFT-optimized structures. We find positive defect formation energies for cells containing Ga located at both doping sites (Figure 2a and Table S1), implying some degree of non-equilibrium growth necessary (i.e., MBE growth) in both cases. However, the formation energy of the substitutional Ga-doped structure (0.314 eV) is markedly lower than the interstitial positions (1.946 eV), indicating that substitutional dopants are expected to be more energetically favorable in the material. This preference is inherently connected to the requirement of the Ge host to undergo an expansion to accommodate the larger Ga atom at the interstitial site, as reported in Table S1. From our DFT calculations, substituting Ge with Ga at these doping levels is expected to impart a relatively small out-of-plane expansion of the lattice parameter of $\sim$0.12%, resulting in a slight tetragonal distortion to the structure (Table S1) to accommodate the heterovalent group-III dopant in a lattice-matched film.
To verify our theoretical calculations of doping site preference in hyperdoped Ge, we perform synchrotron X-ray absorption spectroscopy measurements which have previously demonstrated strong capabilities for determining the lattice position of dopant species Zhang et al. (2024). From our calculations (Figure S6), X-ray absorption at the Ga K-edge is inherently sensitive to the Ga-related local fine structure and can be used to unambiguously resolve its location within the lattice. Experimentally, we employ extended X-ray absorption fine structure (EXAFS) analysis, as measured from the Ga:Ge thin film in a fluorescent mode. X-ray absorption data is shown in Figure 2b measured from various points across the sample, normalized and then averaged together with the $k^{2}$-weighted EXAFS signal shown in the inset. The corresponding analysis of the Fourier transform of this data is presented in Figure 2c. The EXAFS data are best described by a substitutional doping model the strongest Ge-Ga nearest-neighbor scattering pair at 2.447 Å, in excellent agreement with the calculated bond length of 2.439 Å (Figure 2a). To rule out other viable dopant coordination environments, such as $\alpha$-Ga segregation and the interstitial lattice position, we simulate the expected profile for these alternative fine-structure environments, shown in Figure S7. These simulations, in combination with no significant fine structure observed between 2.5 and 3.8 Å in the experimental Fourier transform, we confirm that the Ga dopants preferentially occupy substitutional lattice positions, in agreement with our ab initio calculations. Furthermore, the relatively large Debye–Waller factors ($\sigma^{2}$( Å2)) derived from the EXAFS fits (Table S2) highlight the role of both lattice strain and Si diffusion in the Ga:Ge layers.
Using the optimized structures in Figure 2a with the known dopant arrangement, we now study how the electronic structure is perturbed through large incorporations of heterovalent Ga dopants. While a slight tetragonal distortion is anticipated in strained material due to substrate clamping, we find that the calculated band structure is insensitive to this distortion and we use the high-symmetry cubic Brillouin zone and paths for simplicity. The $k$-resolved band structure is presented in Figure 2d and the $k$-integrated DOS is in Figure S8a. An evaluation of the DOS projected onto different atomic orbitals in Figure S8b shows that new, localized states emerge through the hybridization of both the Ga and Ge p-orbitals. As expected, Ga substitution induces p-type doping in the system, with the Fermi level shifting 1.01 eV into the valence band (see Figure 2d). This corresponds to a hole carrier concentration of $5.5\times 10^{21}$ cm-3 that closely approximates the hole concentration from our Hall measurements. Furthermore, due to the lifting and flattening of the bands near the R point of the cubic Brillouin zone, we predict this system to be a candidate narrow-band superconductor. By plotting the Fermi surface and separating the contribution of each of the three bands crossing the Fermi level (band i, ii, and iii), we observe in the surface for band iii, hole pockets of high-mass carriers that cross the Fermi level at the corners of the Brillouin zone (Figure 2e). This feature is clearly represented in the 2D projected cross-sections of the doped Ga:Ge Fermi surface (Figure S9) and is highly sensitive to the position of the Fermi level. Lowering it by 40 meV to simulate a slight increase in the hole concentration will both enlarge the corner hole pocket projections for band iii and introduce similar features into the projected corners of band ii (Figure S10).
Epitaxial interfaces and phonon structure
Figure 3: Cross-sectional electron microscopy reveals coherent crystalline interfaces. a Schematic of the film structure as grown by MBE. b-c Cross-sectional TEM/EDS imaging of the $T_{\textup{c}}$ = 3.5 K hyperdoped sample presenting the compositional profile across the entire film thickness. d Low-magnification cross-sectional HAADF-STEM image of the film. e Zoom-in on the film/substrate and Ga:Ge/Si interfaces, displaying coherent crystalline interfaces. f Raman back-scattering spectra of undoped (black) and doped (red) structures using 532 nm laser excitation at room temperature. The inset highlights the common emergence of highly anharmonic second-order optical bands, identified as the active LO mode here due to the Raman selection rules for backscattering from a (001) Ge surface. For more details regarding the Raman spectra, please see Figure S12.
Figure 3a presents a schematic structure of the grown film with $T_{\textup{c}}$ = 3.5 K, in which two hyperdoped Ge layers are grown surrounding a thin Si buffer layer. Cross-sectional transmission electron microscopy (TEM) and energy dispersive spectroscopy (EDS) measurements are presented in Figure 3b-c and show the spatially homogeneous elemental profile of the Ga dopants throughout the hyperdoped layer. Notably, no distinct signatures of Ga clustering are observed compared to earlier reports where annealing was used to activate dopants Sardashti et al. (2021). Likewise, these MBE-grown materials also exhibit abrupt interfaces and a smooth surface amenable to future planar device fabrication as opposed to previous MBE studies Strohbeen et al. (2023). A view of the atomic structure using low-magnification high-angle annular dark field scanning transmission electron microscopy (HAADF-STEM) is presented in Figure 3d, with a corresponding expansion of the key film/substrate and Ga:Ge/Si interfaces contained in Figure 3e. To verify domain structure, a wider view of the STEM lamella is presented in Figure S11 where we see no distinct signs of polycrystalline domain formation nor significant surface roughness. Owing to the small amount of mismatch strain and thin layer thicknesses, the growth of Ga:Ge is commensurate with the underlying Ge substrate and the atomic planes on both sides of the interface are in perfect registry. We note that in the current study, we measure a barrier thickness of $\sim$1 nm whereas we estimate a thickness of 5 nm of Si or 10 nm of Ge to be necessary for the demonstration of Josephson tunneling. However, this demonstration of fine thickness control, commensurate interfaces, and homogeneous dopant distribution exhibits the necessary structural and chemical requirements to reduce the disorder in superconductive quantum devices and thus positions these hyperdoped trilayer structures to better mitigate microwave photon loss in quantum information applications Shim and Tahan (2014).
We observe that the heavy p-type nature produces a new, softened mode near 280 cm-1 Compaan et al. (1983) in the (zone center) phonon spectra as seen in Raman backscattering spectroscopy (Figure 3f) compared to the undoped case. Details regarding this measurement are presented in the Methods and in Figure S12. This mode screens the Ge-Ge optical mode, now coupled to the large hole concentration, leading to significant reduction in intensity. Lattice dynamics calculations (Figure S13) of pure and hyperdoped systems support this observation and reveal phonon softening in hyperdoped Ge. Evaluating the Raman spectra in Figure 3f, both doped and undoped samples yield comparable Raman signals related to the Si-Si and Si-Ge-type sublattices. For the Si-Si optical vibration originating from the thin Si layers, the low dimensionality and diffusion of Ge and Ga atoms into this region lead to a significant mode softening and a strong Si-Ge band near 400 cm-1 (Figure 3c) Fournier-Lupien et al. (2013). As both doped and control films exhibit similar Si-Si, Si-Ge, and Ge-Ge bands, we infer high Ga doping does not appreciably affect the growth process. We further compute the Eliashberg spectral function Giustino (2017) (see Figure S14) yielding the total coupling strength, $\lambda=0.41$, from which we calculate $T_{\textup{c}}=0.77$ K Allen and Dynes (1975); McMillan (1968), providing evidence of a conventional phonon-mediated BCS-type pairing mechanism. This coupling primarily originates from the optical phonon modes similar to previous observations in silicon and diamond Ekimov et al. (2004); Bustarret et al. (2006). More details of this calculation are reported in the Methods section.
Dopant-driven crystalline distortion
Figure 4: X-ray scattering measurement of tetragonal crystalline distortion. a 2D GISAXS pattern of the Ga:Ge epitaxial film recorded at the critical angle (0.26) of the superconducting layer. The intensity map is on a linear scale. b The out-of-plane X-ray reflectivity measurement of the Ga:Ge film with a fit made using the REFLEX software package Vignaud and Gibaud (2019). c Cross-section of the electron scattering length density (ESLD) of the epitaxial stack derived for the fitting model of the fringes in b. d 2D GIWAXS pattern of the Ga:Ge thin film record at the critical angle (0.125). The color map is on a log scale. The family of crystallographic planes giving rise to the intense Bragg reflections of the hyperdoped layer are identified with arrows. e Corresponding integrated GIWAXS profile (qxyz) and its subsequent structural refinement (Le Bail method). The refined unit cell parameters are displayed and an expansion of the split 220 family of Bragg peaks is shown in the inset to highlight the reduced symmetry of the clamped Ga:Ge epitaxial layer in comparison to the undoped Ge control.
To resolve the microstructure of the hyperdoped heterostructure we use a combination of synchrotron-based grazing incidence X-ray scattering experiments. The 2D grazing-incidence small-angle x-ray scattering (GISAXS) pattern presented in Figure 4a is recorded at the critical angle of the Ga:Ge layers (0.26), well above the critical angle of the less dense Si layers (Figure S15). Kiessig fringes due to the interference of scattered photons within the Ga:Ge film provide evidence of sharp interfaces between the layers. X-ray reflectometry (XRR) analysis is used to derive the electron density profile and compositional grading parameters (rms roughness) of the film in the normal direction. Based on the XRR fit shown in Figure 4b, compositional grading is confirmed between the layers and diffusion is revealed to be asymmetrical with respect to the stack, with larger roughness values determined for Si grown on top of Ga:Ge, compared to Ga:Ge grown on the Si layer. From the XRR fitting, we observe finite diffusion of Ge into the barrier layer, reducing the epitaxial mismatch strain and helping to preserve epitaxial coherence across this interface. Comparing the XRR electron density profile to the characteristic superconducting length scales as calculated (details in the Supplemental information) from the magnetic field dependance of this sample, presented in Figure S16, we note that the width of the diffusion profile falls well below one coherence length, $\xi$. This is in contrast to the secondary ion-mass spectroscopy (SIMS) profiles in Figure S17, which suggests a long tail of Ga diffused into the Ge substrate. We believe this disagreement to be the result of the sputtering process smearing out the backside interface during our SIMS measurements. Regardless, in the worst case scenario the Ga:Ge/substrate interface is approximately 1 $\xi$ away from the level at which the Ga dopants fall to a concentration below $1\times 10^{17}$ cm-3 (instrumental resolution). Thus, following Anderson’s theorem Anderson (1959) we do not expect significant perturbation of the observed superconducting state due to impurity diffusion since the characteristic length scale of our superconductor, $\xi$, is larger than the characteristic diffusion length scale observed in our experiments.
[…]
Conclusions
Our results establish that superconductivity in MBE-grown Ga:Ge thin films arises from substitutional Ga incorporation rather than Ga-clustering, allowing for the creation of coherent superconductor-semiconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal ($\beta$-Sn) distortion, in line with our DFT results. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level deep into the valence band and flattens the electronic bands near the R point, likely promoting superconductivity. These results show significant promise for future device implementation of superconducting group IV materials; however, challenges yet remain. The band offsets of these superconducting materials interfaced with their semiconducting counterparts and the effect of long-range disorder (large kinetic inductance) in device applications are unable to be fully addressed in the current study. Looking ahead, it is necessary to further develop the proof-of-principle trilayer structures discussed here into full-fledged Josephson junction devices. Furthermore, an equivalent Ga floating layer forms during silicon doping Kesan et al. (1991) allowing for a natural extension of this study towards growth of superconducting silicon. This would enable the trade of relatively lossy germanium substrates for state-of-the-art low loss silicon substrates for use in quantum information devices. The findings presented here highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge thin films and underscore the potential of group IV covalent superconductors as a scalable platform for epitaxial superconductor-semiconductor heterostructures.
Acknowledgments
S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation. P.J. acknowledges fruitful discussions with Ross McKenzie and Ben Powell. P.J.S and J.S. acknowledge funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the Australian National Computational Infrastructure and Pawsey Supercomputing Research Centre through the National Computational Merit Allocation Scheme. This work was partially supported by the Australian Research Council under the following grants: DE230100173 (J.A.S.), LP210200636 (Y.-H.C., P.J.), DE220101147 (C.V.). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University.
Author contributions:
Conceptualization: P.J., J.S., and E.D. Methodology: J.A.S., P.J.S., C.V., S.S.-R. Investigation: J.A.S., P.J.S., C.V., A.B., A.D., Y.-H.C., J.V.D., F.K., A.L., D.P., S.S.-R. Visualization: J.A.S., P.J.S. Funding acquisition: J.A.S., P.J., J.S. Project administration: J.A.S., P.J.S., P.J., J.S. Supervision: J.A.S., P.J., J.S., E.D., L.W. Writing – original draft: P.J., J.A.S., P.J.S. Writing – review and editing: J.A.S., P.J.S., C.V., P.J., J.S.
Competing interests:
There are no competing interests to declare.
Methods
[…]
Hall bar device fabrication and measurement. A $\times5$ chip is first cleaned in acetone for 2 min, followed by isopropyl alcohol for 1 min, and then blow dried with compressed air. The chip is then dehydrated on a hotplate in a fume hood at 110C for 1 minute and then immediately placed into a spin coater. AZ 5214 photoresist (PR) is spun onto the chip at 4000 rpm for 60 s. Edge beads are removed using acetone and then the chip is placed back onto the hotplate at 110C for 1 min for the PR post-bake. The photolithography was conducted with a Süss MJB3 mask aligner using hard contact. The PR was exposed for 9 s, developed for 30 s in AZ 1:1 developer, and then rinsed in DI water for 15 s. The Hall bar was etched in an Oxford PlasmaPro System 100 Cobra inductively-coupled plasma (ICP) dry etcher system using a chlorine gas chemistry of 5 sccm Cl2, 5 sccm BCl3, 5 sccm Ar for 30 s at an ICP power of 500 W and a 75 W bias. This etch results in a step height of 140 nm as measured in an atomic force microscope, confirming a deep etch into the Ge substrate. The resist is stripped in AZ 400T PR stripper on a hotplate at 80C for 20 minutes, cleaned in isopropyl alcohol, and then blow dried. The device is wire bonded to a custom daughterboard and then cooled down in an Oxford Triton dilution refrigerator with a base temperature of 15 mK and a 14 T single-axis magnet.
Four-point transport measurements. Electrical characterization is done in an Oxford TeslatronPT He4 cryostat using a HelioxVT He3 probe insert with a base temperature of $\sim$ 270 mK (HelioxVT) and magnetic field capabilities up to 12 T. Measurements are collected in a similar fashion to Ref. Strohbeen et al. (2023), using a standard Van der Pauw wiring configuration on square pieces cleaved from near the center of each wafer. On-chip contacts are made via annealed In-Sn eutectic alloy at each of the four corners which are then wired to a daughterboard using gold wire.
First-principles calculations. First-principles calculations were performed using DFT as implemented in VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996), except for the electron-phonon coupling calculations where we employed Quantum ESPRESSO Giannozzi et al. (2017). In VASP, projector augmented wave (PAW) pseudopotentials including semicore 3d electrons were adopted, with a plane-wave cutoff of 550 eV. A $7\times 7\times 7$ $\Gamma$-centered k-point mesh was used to sample the Brillouin zone of the conventional 8-atom unit cell of Ge (cubic diamond structure, space group $Fd\overline{3}m$). Structural optimizations were performed using the PBEsol exchange-correlation functional Perdew et al. (2008), which yields lattice parameters in excellent agreement with the experiment (see Table S1), and a convergence threshold of 0.005 eV/Å on all forces.
For the Ga-doped Ge structure with substitutional fine-structure, one host atom in the 8-atom unit cell is replaced by a Ga atom, corresponding to a doping concentration of 12.5%. Conversely, interstitial doping is modeled by placing one Ga atom at interstitial doping sites, representing a doping concentration of 11.1%. To model the effect of substrate clamping in Ga-doped Ge, the in-plane lattice parameters were fixed to the pristine Ge values, and only the out-of-plane lattice parameter was allowed to relax. K-edge spectra were simulated using the supercell core-hole method Karsai et al. (2018) within a $2\times 2\times 2$ supercell containing 64 atoms (72 atoms in the case of interstitial doping). The formation energy of a neutral Ga dopant in the Ge crystal is calculated as $E_{\textup{Form}}=E_{\textup{d}}-E_{\textup{p}}-\sum_{i}n_{i}\mu_{i}$, where $E_{\textup{d}}$ is the total energy of the structure containing the defect, $E_{\textup{p}}$ is the total energy of the pristine structure, $n_{i}$ is the number of atoms of species $i$ being added ($n>0$) or removed ($n<0$), and $\mu_{i}$ is the chemical potential. For Ga, this is the energy per atom of a Ga crystal (base-centered orthorhombic structure).
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional Tran and Blaha (2009) based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge Rödl et al. (2019); Fadaly et al. (2020), which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included. A denser $20\times 20\times 20$ $k$-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a $31\times 31\times 31$ $k$-mesh was used. Vibrational properties (shown in Fig. S15) were computed using finite differences Togo and Tanaka (2015) within supercells containing 216 atoms ($3\times 3\times 3$ supercells of the 8-atom unit cell), employing $3\times 3\times 3$ $k$-point meshes and displacement amplitudes of 0.015 Å.
We used Quantum ESPRESSO Giannozzi et al. (2017) to perform the electron-phonon coupling calculations, using norm-conserving pseudopotentials from the PseudoDojo library v0.4.1 van Setten et al. (2018) with standard accuracy and the PBEsol exchange-correlation functional, and a plane-wave cutoff of 80 Ry (1088 eV). We adopted the same model of substitutionally Ga-doped Ge, ignoring for simplicity the effect of substrate clamping. Relaxing the unit cell after Ga substitution results in a small lattice expansion of 0.09%. Spin-orbit coupling was neglected. We verified that the resulting Fermi level shift into the valence bands (0.96 eV) is consistent with the one obtained with the MBJLDA functional within VASP when spin-orbit coupling is neglected (0.93 eV). We calculated the Eliashberg spectral function $\alpha^{2}F(\omega)$ defined as Giustino (2017):
$\alpha^{2}F(\omega)=\frac{1}{N_{\textup{F}}N_{\mathbf{k}}N_{\mathbf{q}}}\sum_{mn\nu,\mathbf{k}\mathbf{q}}|g_{mn\nu}(\mathbf{k},\mathbf{q})|^{2}\,\delta(\varepsilon_{n\mathbf{k}}-\varepsilon_{\textup{F}})\delta(\varepsilon_{m\mathbf{k}+\mathbf{q}}-\varepsilon_{\textup{F}})\delta(\hbar\omega-\hbar\omega_{\mathbf{q}\nu}),$ (1)
where $N_{\textup{F}}$ is the density of states at the Fermi level, $N_{\mathbf{k}}$ and $N_{\mathbf{q}}$ are the total number of $\mathbf{k}$ and $\mathbf{q}$ points used to sample the Brillouin zone, $\varepsilon_{n\mathbf{k}}$ is the electron energy with band index $n$, $\varepsilon_{\textup{F}}$ is the Fermi energy, and $\hbar\omega_{\mathbf{q}\nu}$ is a phonon energy with momentum $\mathbf{q}$ and branch $\nu$. $g_{mn\nu}(\mathbf{k},\mathbf{q})$ are the electron-phonon matrix elements Giustino (2017) calculated using density-functional perturbation theory. The total electron-phonon coupling strength $\lambda$ is then obtained as $\lambda=2\int d\omega\frac{\alpha^{2}F(\omega)}{\omega}$. Using the semi-empirical Allen-Dynes-McMillan equation, the superconducting critical temperature $T_{\textup{c}}$ is estimated as McMillan (1968); Allen and Dynes (1975):
$k_{\textup{B}}T_{\textup{c}}=\frac{\hbar\omega_{\textup{log}}}{1.2}\,\mbox{exp}\left[\frac{-1.04(1+\lambda)}{\lambda-\mu^{\ast}(1+0.62\lambda)}\right],$ (2)
where $k_{\textup{B}}$ is Boltzmann’s constant, $\omega_{\textup{log}}$ is the logarithmic average of the phonon frequencies, given by $\omega_{\textup{log}}=\mbox{exp}\left[\frac{2}{\lambda}\int_{0}^{\infty}d\omega\frac{\alpha^{2}F(\omega)}{\omega}\log\omega\right]\!,$ and $\mu^{\ast}$ is the Coulomb pseudopotential, here set to 0.1 Allen and Dynes (1975). To converge the Brillouin-zone integrals in Eq. (1), we employed a $20\times 20\times 20$ and $5\times 5\times 5$ $k$ and $q$-mesh, respectively, and a Gaussian broadening of 0.04 Ry for the electronic double-delta functions.
Scanning transmission electron microscopy. Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64-200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 $\times$ 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy. Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal micro-Raman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10x, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 cm-1), instrument calibration was verified through checking the position of the Si band at $\pm$520.7 cm-1, and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S12). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
Synchrotron-based GISAXS/GIWAXS. To identify the crystalline phases and microstructure of films, synchrotron-based grazing incidence wide-angle X-ray scattering (GIWAXS) data were collected at the small angle X-ray scattering/wide angle X-ray scattering (SAXS/WAXS) beamline at the Australian Synchrotron Kirby et al. (2013): 2D scattering patterns were recorded using a wavelength of either 1.340375 Å (9.25 keV) or 0.619924 Å (20 keV), using an energy-selective Pilatus 2M detector. The measurements made using a 20 keV beam energy were to increase the observed Q-range for the structural refinements, using an energy threshold of 15 keV to suppress the X-ray fluorescence signals arising from the excited K-edges of Ge (11.1 keV) and Ga (10.37 keV). Experiments at 9.25 keV had no such issues. Two different sample-to-detector distances were employed to record both SAXS (5.5 m) and WAXS (0.45 m) profiles, each calibrated using a silver behenate reference standard. The sample and detector were enclosed in a vacuum chamber to suppress air scatter. Scattering patterns were measured as a function of the angle of incidence, with data shown acquired with an angle of incidence near the critical angle (for a given X-ray energy) to maximize scattering intensity from the sample. All collected 2D images were azimuthally integrated using PyFAI Ashiotis et al. (2015) and processed using a custom Python routine. The resulting unit cell models were refined using the La Bail method implemented in Fullprof Rodríguez-Carvajal (1993), a comprehensive analysis software. In line with our other characterization techniques, our GIWAXS experiments failed to detect Ga droplet formation or segregation within the film. Due to the relatively small scattering volume of the Si layers, no additional peaks are indexed to this phase either.
Synchrotron-based EXAFS experiments. Ga K-edge X-ray absorption spectra (XAS) were captured at the Australian Synchrotron (ANSTO in Clayton, Victoria) MEX1 beamline in a similar fashion to previously reported experiments Zhang et al. (2024). A double-crystal Si (111) monochromator equipped with focusing optics was used to reduce harmonic content while producing excitation energy. An inline Ga2O3 reference (the initial peak of the first derivative occurs at 10.3687 keV) was used to calibrate the monochromator at Ga-K absorption edge with E0 set to its known value of 10.3671 eV. All acquisitions were performed on thin film materials in a fluorescent mode. Brief XAS scans of specific samples were conducted to verify the stability of the materials under at least several minutes of X-ray exposure before the full acquisitions. Samples were measured with varying energy intervals over the pre-edge (5 eV) and the XANES region (0.25), with 0.035 intervals in $k$-space over the EXAFS. At low-$k$ an integration time of 1 s was employed per interval, with longer integration times weighted toward high-$k$ portions of the spectrum, up to a maximum value of 9 in $k$-space (max of 2 s). Multiple scans were collected at different positions of the sample. Data processing including background subtraction, scan averaging, edge-height normalization, and rebinning was performed based on the Athena program Ravel and Newville (2005). For normalization of energy spectra and removal of background, the pre-edge range was set between -150 eV and -30 eV, while the normalization range spanned from 80 to 300 eV post-edge. The order of normalization was designated as 3. For the presented EXAFS data, the $k$-weight was configured as 2, and the $k$-range for the forward Fourier transform was defined from 0.8 to 8.5.
EXAFS fine-structure modeling. Fits were conducted in the ARTEMIS, part of the IFEFFIT software package Ravel and Newville (2005) using a similar procedure to previous studies Zhang et al. (2024). The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package Ravel and Newville (2005), with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table S2). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.
[…]
Supplemental Information
S1 State of hyperdoping in germanium
Figure S1: a Comparison of this study against literature reports for superconducting $T_{\textup{c}}$ as a function of Hall carriers in Ga-doped germanium. The colored star, square, and triangular points were prepared via ion implantation Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021) while the black circles (this study) and black diamond Strohbeen et al. (2023) are MBE-grown samples. b Representative AFM image of the surface of the sample circled in red in a. We observe an RMS roughness of roughly 1.65 nm, a significant improvement over previous MBE reports Strohbeen et al. (2023). The inset presents the normalized height distribution, $\rho$, across the entire 15 $\mu m\times$15 $\mu m$ image.
S2 Material optimization
Figure S2: RHEED line cuts for superconducting hyperdoped Ge thin film. a Ge substrate $\langle 110\rangle$-type direction, taken at room temperature just before deposition of Ga:Ge film. b Initial surface of Ga:Ge film. c RHEED pattern after the first 10nm of Ga:Ge deposited. d After the Si spacer layer and initial growth of the second Ga:Ge layer. e End of Ga:Ge growth, before Si cap is deposited. We note a superimposed amorphous ring over a faint $\langle 110\rangle$-type direction diffraction pattern. This is consistent with previous observations that Ga-adsorption at low temperatures creates a floating layer at the surface from which the incorporated dopants are taken Kesan et al. (1991).
Film growth is monitored using reflection high-energy electron diffraction (RHEED) along a $\langle 110\rangle$ cut. Representative RHEED images throughout the growth steps are presented in Figure S2. The haziness observed in Figure S2b-e is caused by a surface floating layer of Ga that forms during the growth process Kesan et al. (1991). The Ga shutter is closed during Si deposition, however, we note that the floating layer slightly incorporates into the Si layer as indicated by the increased intensity in the specular RHEED reflection in Figure S2d compared to Figure S2c. To reduce parasitic heating of the substrate, the hot Ga crucible was backed up within the vacuum chamber by 4 in to reduce radiative heating of the growth surface when the shutter is open.
Figure S3: Superconducting $T_{\textup{c}}$ comparing the effects of silicon as a function of hole concentration and growth temperature. The circular data points are the samples where we replaced the Si tunnel barrier and capping layers with germanium while the triangular points are the samples that contain silicon. The black points are samples in which just a single layer of hyperdoped germanium is grown and the red points are the samples where we grew bilayer heterostructures. We observe a tendency for Si-containing samples to exhibit marginally increased carrier concentration and superconducting $T_{\textup{c}}$. A trend in $T_{\textup{c}}$ as a function of growth temperature is further demonstrated where we optimize the superconducting transition temperature at growth temperatures of 95-150C.
The scatter in $T_{\textup{c}}$ observed in Figure S3 for samples grown at similar conditions is attributed to small variations in atomic fluxes caused by drift during the course of film growth as well as the slight turbulence present within the Ge melt Lee et al. (1998). In principle, changing the Ge deposition source from an electron beam deposition source to a different technology would help mitigate these variations.
Figure S4: Hole concentration as a function of substrate temperature during growth. The same data point identification in Figure S3 is also used here. a Presents the entire set of samples, where we observe two outlier samples: one in which carrier concentration is drastically reduced compared to the other samples and one with significantly higher carrier concentration. We speculate poor dopant activation and Ga droplet formation as the cause of these two samples, respectively. We zoom in on the rest of the samples in b, where we see that a general downward trend is observed as substrate temperature increases. This trend is consistent with an increased growth temperature promoting Ga segregation, reducing the carrier concentration of the films.
Figure S5: Hall bar device fabrication and transport characterization of a bilayer heterostructure. a Presents an optical image of the fabricated device after wirebonding. We measure under current bias using a lock-in amplifier at low frequency to reduce noise in our measurements. b Shows the longitudinal resistance data upon cooldown of the device. These measurements are taken in a dilution refrigerator to ensure the sample remains superconducting down to mK temperatures with no re-entrant phase. However, we note that dilution fridges are non-ideal for measurements at higher temperatures where we see these vertical line artifacts during our measurement. This is caused by a delay in the thermometry of our fridge in updating the temperature of the mixing chamber plate. The transition temperature exists around 3 K, similar to the Van der Pauw measurement reported in Figure 1c on the same wafer. c Longitudinal resistance as a function of out-of-plane magnetic field also reports a critical magnetic field of the superconducting state to be $\sim$1 T, similar to what is reported for the Van der Pauw device presented in Figure S16a. d Hall trace on this device from which we calculate the carrier concentration to be $4.67\times 10^{21}$ cm-3, which is comparable to the Van der Pauw measurement of $4.15\times 10^{21}$ cm-3. From this data we conclude that the hyperdoped germanium thin film material is highly robust to typical photolithography and chlorine dry etch processing (more details are found in the Methods section). The inset schematic cartoons in b, c, and d represent the specific pins being measured for the presented traces based on the optical image shown in a. The dark boxes represent the pins being measured in these schematics.
S3 X-ray absorption measurement alignment and fitting
Figure S6: Calculated X-ray absorption spectra of Ga K-edge for Ga:Ge crystals with substitutional and interstitial doping sites.
Figure S7: Comparison of the simulated fine structure derived from the CIF files of Ga:Ge interstitial system (DFT calculated system in Fig. 2a of main text) and amorphous Ga metal Sharma and Donohue (1962). Note that these Fourier transforms have been phase corrected, for simplicity. The arrows highlight the relatively large difference in experimental bond length of the nearest-neighbor scattering pair and those derived in the simulations.
S4 Impact of dopants on electronic structure
Figure S8: a Total density of states (DOS) calculated for the pure Ge crystal and the hyperdoped Ga:Ge shown in Fig. 2a of the main text, with Ga at interstitial (int.) and substitutional (sub.) lattice positions. The energies are aligned to the valence band maximum (VBM), set as the zero of the energy. The Fermi level in each case is indicated by a dashed vertical line. We observe further evidence that interstitial doping is unfavorable for for this system as it destroys the “clean” electronic bands of the host Ge semiconductor as evidenced by the disappearance of the parent Ge band gap and the emergence of metallic states. In contrast, substitutional Ga doping is expected to form acceptor states below the valence band edge, while the overall band structure near the gap remains largely unperturbed (see Figure 2d). b Total and projected DOS of the hyperdoped Ga:Ge crystal with substitutional fine-structure.
Figure S9: a 3D Fermi surface for substitutionally doped Ga:Ge, showing the (011) plane. b 2D Fermi surface cut on the (011) plane, showing the contribution of bands i, ii, and iii. Hole pockets from band iii are visible in the R-point corners.
Figure S10: Fermi surface obtained after shifting the Fermi level by 40 meV below the value shown in Fig. 2d of the main text. a Contribution of bands i, ii, and iii to the 3D Fermi surface, with the color corresponding to the Fermi velocity. b (011) plane crossing the 3D Fermi surface. c 2D cut on this plane, showing the contribution of each band. The predicted narrow- or flat-band conditions are shown to be sensitive to the position of the Fermi level, as the pockets of high-mass carriers from bands $i$ and iii at the corners of the Brillouin Zone (R point) are shown to enlarge with a relatively minor shift in the Fermi level.
S5 Domain structure and near-surface Raman spectroscopy
Figure S11: a Low magnification image of the STEM lamella presented in Fig. 3 in the main text. We observe no domain structure as reported in previous reports of MBE-grown hyperdoped Ge films Strohbeen et al. (2023). The red dashed-line box is the region of higher magnification presented in b and Figure 3d.
Figure S12: a Raman Scattering penetration depth ($\mathrm{d_{\mathrm{RS}}}$) as a function of the laser excitation energy, determined by considering the absorption coefficients Aspnes and Studna (1983) for both the incoming ($\alpha_{\mathrm{out}}$) and the outgoing ($\alpha_{\mathrm{in}}$) light: $\mathrm{d_{\mathrm{RS}}}=1/(\alpha_{\mathrm{in}}+\alpha_{\mathrm{out}})\approx 1/\alpha_{in}$. b Comparison of Raman spectra recorded from Ga:Ge and Ge control MBE films at different excitation energies. At longer excitation wavelengths (785 nm), both the control and target films exhibit a comparable Raman response. Probing the surface-sensitive information with shorter wavelengths, the differences between the vibrational structure of the films become clearer. The appearance of disorder-activated LA modes below 200 cm-1 (from both Ge-type and Si-type sublattices) in the scattering volume yields comparable scattering intensities under these conditions for the control and doped sample. We suggest the origins of these bands are a combination of (local) lattice strain, phonon confinement and (partial) satisfaction of the resonance conditions of the Ge crystal $E_{1}+\Delta_{1}$ gap Alonso and Cardona (1988). Together, these features break the normal Raman scattering selection rules of a normal (001)-oriented bulk Ge film, i.e., resembling more the Raman spectra recorded using 785 nm laser light.
S6 Calculation of electron-phonon coupling
Figure S13: Calculated first-order phonon DOS for pure Ge and hyperdoped Ga:Ge crystal.
Figure S14: Eliashberg spectral function $\alpha^{2}F(\omega)$ calculated for substitutionally doped Ga:Ge and cumulative electron-phonon coupling strength $\lambda(\omega)$, showing that the coupling originates mostly from optical phonon modes. The total integrated $\lambda$ is 0.41, yielding a value of $T_{\textup{c}}=0.77$ K from the McMillan-Allen-Dynes formula [Eq. (2)].
To investigate the role of phonons as an underlying mechanism for the observed superconducting state, we perform additional calculations on the electron-phonon coupling strength within our model of hyperdoped Ge. Specifically, we compute the Eliashberg spectral function $\alpha^{2}F$ Giustino (2017), details of which are reported in the Methods section, to yield the total coupling strength, $\lambda$. From these quantities, the superconducting transition temperature, $T_{\textup{c}}$, is estimated using the McMillan-Allen-Dynes formula Allen and Dynes (1975); McMillan (1968). We calculate $\lambda=0.41$ and $T_{\textup{c}}=0.77$ K, providing evidence of a conventional phonon-mediated BCS-type pairing mechanism within this system. The Eliashberg spectral function reported in Figure S14 shows that this coupling originates primarily from the optical phonon modes, similar to what has been previously observed in the case of boron-doped diamond Ekimov et al. (2004) and silicon Bustarret et al. (2006).
S7 Synchrotron penetration depth and calculation of characteristic superconductor length scales
Figure S15: The calculated penetration depth of the different material layers at the synchrotron energies used for grazing incidence experiments. For the range of angles below and at the critical angle, the X-ray electromagnetic field only interacts a short distance below the film surface due to the evanescent damping (5 - 10 nm), and constitutes the defined evanescent regime.
Figure S16: Out-of-plane a and in-plane b critical magnetic field behavior for the Ga:Ge sandwich heterostructure presented in Figure 1c and Figure 3.
Figure S17: Secondary-Ion Mass Specroscopy data for two samples investigated in this study, a PJS146 ($T_{\textup{c}}\sim$ 2.5 K) and b PJS147 ($T_{\textup{c}}\sim$ 3.5 K).
The magnetic field dependence of longitudinal resistance within the Josephson structure is presented in Figure S16a and b for out-of-plane and in-plane field configurations, respectively. At 300 mK we observe the out-of-plane $H_{\textup{c}}$ to be 0.79 T, decreasing to 0.14 T at 2 K. In-plane $H_{\textup{c}}$ is found to be 2.51 T, decreasing to 0.78 T at 2 K. We calculate the coherence length from this data using the following equation:
$\xi=\sqrt{\frac{\Phi_{\textup{0}}}{2\pi H_{\textup{c2}}(0)}}$ (3)
where $\Phi_{\textup{0}}$ is the magnetic flux quantum and $H_{\textup{c2}}(0)$ is the upper critical magnetic field extrapolated to zero Kelvin. London penetration depth is also calculated using:
$\lambda_{\textup{L}}=\sqrt{\frac{m^{*}}{2\pi n_{\textup{s}}q^{2}}}$ (4)
where $m^{*}$ is the effective mass of the hole carriers which we take to be the weighted average, $0.28m_{\textup{e}}$, and $n$ is the superfluid density which we obtain from our Hall measurements as $2n_{\textup{s}}=n_{\textup{h}}$. We calculate the penetration depth ($\lambda$) and coherence length ($\xi$) on the trilayer sample presented in Figures 3 and 4 to be 29 nm and 13.4 nm, respectively, similar to what has been previously observed in boron-doped diamond superconductors Bustarret et al. (2004).
S8 GIWAXS alignment, control, and stress analysis
[…]
Figure S19: Integrated GIWAXS profile (qxyz) and structural refinement (Le Bail method) of the pure Ge epitaxial layer. Unstrained Ge crystal belongs to the high-symmetry cubic space group $Fd\overline{3}m$ (space group No. 227), characterized by equivalent lattice parameters ($a=b=c$) and a structural refinement of our undoped Si-Ge film aligns well with the reported bulk, cubic value ($a=5.657906\pm 0.0000009$ Å, Vol. = 181.1203 Å3Baker and Hart (1975).
Stress Analysis
The lattice mismatch is accommodated by an elastic strain in the hyperdoped epilayer, giving a biaxial stress of:
$\sigma=2\mu f\cdot(1+\nu)/(1-\nu)$ (5)
where $\mu$ is the shear modulus, $\nu$ Poisson’s ratio and $f$ the misfit parameter. Assuming the hyperdoped Ga:Ge epilayer forms a cubic unit cell when unconstrained, elastic isotropy is assumed in the relevant deformation directions and the misfit parameter is given by:
$f=\frac{(a_{\textup{Ga:Ge}}-a_{\textup{Ge}})}{a_{\textup{Ge}}}$ (6)
where $a_{\textup{Ga:Ge}}$ and $a_{\textup{Ge}}$ are the lattice parameters of the relaxed epilayer and substrate, respectively. We do not have such experimental data for the relaxed hyperdoped unit cell, however, since the epilayers studied here are pseudomorphically grown (Figure 3e), the values obtained from the structural refinement can be related to the relaxed lattice parameter of the Ga:Ge layer through:
$a_{\textup{Ga:Ge}}=a_{\textup{Ge}}[1+P(\Delta d/d)_{\perp}]$ (7)
where $P$ is the elastic parameter of the films McSkimin and Andreatch Jr. (1963). Regardless of the growth direction, the in-plane and out-of-plane biaxial strain components are given by:
$\epsilon_{\parallel}=(a_{\textup{Ga:Ge}}-a_{\textup{Ge}})/a_{\textup{Ge}}$ (8)
$\epsilon_{\perp}=\epsilon_{\parallel}(1-1/P)$ (9)
respectively. For our thin films grown on Ge (100); $P_{100}=C_{11}/(C_{11}+2C_{12})=0.5709$. To implement our calculation, we approximate the elastic constants $C_{ij}$ of the hyperdoped film to be comparable to the pure Ge crystal McSkimin and Andreatch Jr. (1963) (as there are no relevant data for GaGe alloys). Based on our hole carrier concentrations, the Ga concentration for the sample studied in Figure 3e is $\sim$11.22% and has a calculated relaxed (cubic) lattice parameter of $a_{\textup{Ga:Ge}}$ = 5.6614 Å, which is 0.11% larger than the Ge crystal.
S9 DFT and EXAFS data tables
Table S1: Structural parameters of DFT optimized structures and defect formation energies $E_{\textup{Form}}$. Values are calculated for pure and Ga-doped Ge (both substitutional and interstitial doping).
Structure Lattice Parameter (Å) Volume (Å3) Crystal Structure $E_{\textup{Form}}$ (eV)
a=b, c
Pure Ge 5.674, 5.674 182.64 Cubic
Sub. Ga:Ge 5.674, 5.681 182.87 Tetragonal 0.314
Int. Ga:Ge 5.674, 6.220 200.128 Tetragonal 1.946
Table S2: EXAFS fitting parameters. Values are extracted from fits made to spectra recorded at 295 K using a Ga substitutional model. The lower ranked paths (in italics) are omitted from the actual fit model to avoid overfitting. The errors provided in the parentheses align with the last significant figures of the variable fit values. Further fitting details can be found in the Methods section. The path type can be tracked as: $S$ - single scattering pair; DS - double scattering; R - rattle.
[…]
Ga-Ge (SS) 4 100 2.451(1) 0.0051(4)
Ga-Ge (SS) 12 90.82 4.002(5) 0.0063(2)
Ga-Ge-Ge (DS) 12 5.71 4.450
Ga-Ge-Ge (DS) 24 21.4 4.452(2) 0.010(6) -0.195 0.68(11) 0.015
Ga-Ge (SS) 12 59.45 4.693(2) 0.0087(5)
Ga-Ge-Ge (R) 4 4.09 4.900
Data and materials availability:
Raw data available on Zenodo: doi.org/10.5281/zenodo.17065133 Strohbeen (2025)